OpenStudy (anonymous):

What is the slope of the line that passes through (-2, 2) and (-4, 5)? -3/2 -2/3 3/2 2/3

OpenStudy (anonymous):

slope=y2-y1/x2-x1 =5-2/-4+2 =3/-2 =-3/2